OpenStudy (anonymous):

What is the equation given the points (-5,-3) and (10,0)?

OpenStudy (anonymous):

rise over run. Tell me what that is. -5 is x1 , -3 is y1, 10 is x2 , 0 is y2 y2-y1/x2-x1 What is it.

OpenStudy (anonymous):

0-(-3) / 10 - (-5) Correct?

OpenStudy (anonymous):

Is there another equation i can use?

OpenStudy (aripotta):

0-(-3) / 10 - (-5) was correct for the slope, but you have to simplify further

OpenStudy (aripotta):

what's 0 - (-3)?

OpenStudy (anonymous):

is the final equation y=1/5x-4

OpenStudy (anonymous):

3

OpenStudy (aripotta):

you were partially right the first time. it's y = -1/5x - 4

OpenStudy (aripotta):

oh wait..lemme see here again..

OpenStudy (anonymous):

rise over run gives you your slope. Then .... use the following point slope form: y - yinitial = m(x - xinitial) yinitial = any one of your y points xinitial = corresponding x to above m = slope you found out rearrange to isolate y on one side for y=mx + b

OpenStudy (anonymous):

no it's not negative -1/5x because on the worksheet it says the slope is 1/5x

OpenStudy (aripotta):

yea, sorry. i got mixed up on that :) you were right :)

OpenStudy (anonymous):

oh okay, thank you :)